OpenStudy (luigi0210):

Differentiate implicitly:

OpenStudy (luigi0210):

\[\LARGE tan(x-y)=\frac{y}{(4+x^2)}\]

OpenStudy (anonymous):

for a person at 99 i think you can help yourself XD jk jk

OpenStudy (shamil98):

the tan(x-y) is throwing me off

OpenStudy (shamil98):

and you already know how to do this, so stop asking dumb questions -_-

OpenStudy (luigi0210):

So it turns to.. \[\LARGE \frac{tanx-tany}{1+tanxtany}=\frac{y}{4+x^2}\] So quotient rule on both sides? >.<

zepdrix (zepdrix):

No quotient on the left :o Just derivative of tan, then chain rule.

OpenStudy (luigi0210):

Hmm, you sure we don't use the trig add/sub formulas here? @zepdrix

zepdrix (zepdrix):

I'm not sure why you would want to do that :o Makes a lot more work for you.

OpenStudy (luigi0210):

Hmm, okay you're right xD \[\LARGE sec^2(x+y)*1+y'=\frac{(4+x^2)(y')-(2x)(y)}{(4+x^2)^2}\]?

zepdrix (zepdrix):

Woops, dont' forget the brackets on the chain rule!\[\Large\bf\sf \frac{d}{dx}\tan(x-y)\quad=\quad \sec^2(x-y)\cdot(1-y')\]

OpenStudy (luigi0210):

Right! Question being how do we isolate y' now? xD

zepdrix (zepdrix):

Crap my LoL game is starting +_+ gotta go!!

OpenStudy (luigi0210):

Okay T_T

OpenStudy (luigi0210):

@whpalmer4 ? >.<

OpenStudy (shamil98):

OMG HE PLAYS LEAGUE

OpenStudy (shamil98):

yeaaaaaaaaaaaaaaaaaaaaaaaaaa

OpenStudy (the_fizicx99):

Sham I have you blocked on LoL :>

OpenStudy (shamil98):

i have you blocked on here, and u dont have LoL, and you don't block ppl on LoL either you just put on ignore list. noob

OpenStudy (the_fizicx99):

There is a way ;) beyond ignore list *points* I do, barely enough to consider myself one though :/

zepdrix (zepdrix):

Figure this out Luigi? c: